OpenStudy (anonymous):

Which estimate is most reasonable? Michaela purchases a box of cereal for $3.95, chicken for $5.99, orange juice for $2.10, and a loaf of bread for $1.05. She has $15 with her. About how much change should she expect to receive? $1 $2 $4 $3 1 points

OpenStudy (boonerileycat):

$1 I think

OpenStudy (studygurl14):

15 - 3.95 - 5.99 - 2.10 - 1.05 = ?

OpenStudy (anonymous):

2

OpenStudy (studygurl14):

If you round: 15 - 4 - 6 - 2 - 1 = 11 - 6 - 2 - 1 = 5 - 2 - 1 = 3 - 1 = 2